# ADA-DPM: A Neural Descriptors-based Adaptive Noise Filtering Strategy for SLAM

### v2

@@ 3.3 ADA-DPM for Point Cloud Registration
Importance Scoring Head: The Importance Scoring Head adopts an architecture of stacked MLP and KAN Linear Layer [19], as shown in Figure 3(b). The Importance Scoring Head treats the preliminary estimation of the rigid transformation between $[\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $[\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$ as a regression task. By decoding $[\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $[\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$, it predicts the rigid transformation between them and scores the matching situation as $\sigma_{p}\in\mathbb{R}^{N\times 1}$. During decoding, considering that the rigid transformation between the source and target point clouds satisfies strict mathematical constraints, we use the KAN Linear Layer as the decoder (its mathematical representation[4] is shown Equations (6) and (7)), which overcomes the loss of high-frequency information caused by discontinuous activation functions in traditional MLPs[19].
@@
[…]
It should be noted that directly predicting the rigid transformation $\mathcal{T}\in SE(3)$ between $\tilde{\mathcal{P}}_{src}$ and $\tilde{\mathcal{P}}_{dst}$ requires 12 variables to be predicted. To reduce the optimization difficulty, we adopt a simplification strategy: only predict the rotation transformation, which is represented using the quaternion $q=[s,(x,y,z)]$ (where $s$ denotes the real part of $q$, and $x,y,z$ are the imaginary parts of $q$); the translation transformation is derived through geometric constraints Equation (8). Specifically once $q$ is successfully predicted, we use $q$ to align the $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$. At this point, the difference in the centroid coordinates of the two aligned sets corresponds to the translation vector. In addition, by supervising $q$, implicit supervision can be imposed on the feature extraction network, enabling it to learn to extract feature points that contribute more to the point cloud registration task.
@@
To ensure the performance of the Importance Scoring Head, during the training phase, we specifically design the importance scoring loss $\mathcal{L}_{s}$ to supervise the prediction of $q$ and $\sigma_{p}$. $\mathcal{L}_{s}$ consists of the global reconstruction loss weighted by importance score $\mathcal{L}_{q}$ and the scoring loss $\mathcal{L}_{\sigma}$, where $\mathcal{L}_{q}$ is used to supervise the prediction of $q$, and $\mathcal{L}_{\sigma}$ is used to supervise the prediction of $\sigma_{p}$. Specifically, $\mathcal{L}_{q}$ is represented by the global reconstruction loss weighted by $\sigma_{p}$; $\mathcal{L}_{\sigma}$ is represented by cross-entropy with labels all being 1. Since the network weights are optimized in the direction that reduces $\mathcal{L}_{s}$ during training, for correctly paired point pairs, $\mathcal{L}_{q}$ can be reduced by predicting the correct $q$; for incorrectly paired point pairs, $\mathcal{L}_{q}$ needs to be reduced by predicting a smaller $\sigma_{p}$. Meanwhile, to prevent the model from reducing $\mathcal{L}_{q}$ by predicting small (or even all-zero) $\sigma_{p}$ for all point pairs, $\mathcal{L}_{\sigma}$ calculates the loss with a vector of all ones. This ensures that for correctly paired point pairs, the model predicts $\sigma_{p}$ as close to 1 as possible, and $\mathcal{L}_{q}$ can only be reduced by predicting the correct global rigid transformation, thus avoiding the trivial solution where $\sigma_{p}$ is extremely small or all zero in this case. At this point, $\sigma_{p}$ can be used to represent the contribution of point pairs to registration.
@@ 3.4 ADA-DPM for Loop-Closure Detection
For the Encoder part, we adopt the structure of PointNeXt [38] combined with CLI-GCN. Existing point cloud feature extraction methods (such as PointNet++ [27], VoxelNet [40], and PointNeXt [38] used in this paper) aggregate local features through downsampling. As a result, although the finally extracted compressed feature descriptors have rich semantic features, they suffer from issues such as a single receptive field and low feature resolution. For points in overlapping regions, changes in the size of their neighborhood structures do not affect their similarity; however, for points at the edges of overlapping regions, an increase in the receptive field will cause a sharp decrease in the similarity of their neighborhood structures, which may lead to incorrect loop-closure predictions and introduce errors into robot localization and map construction. To make up for this defect of PointNeXt, we construct the CLI-GCN module to perceive cross-layer multi-scale neighborhood structures, so as to capture richer geometric features.
The CLI-GCN structure is shown in Figure 4, mainly consisting of two parts: (a) Cross-Layer Feature Point Sample: this part searches for neighborhood points for deep feature points within the shallow feature point set (performing neighborhood sampling based on distance), constructing multi-scale neighborhood correlations between deep feature points and shallow associated feature points to enhance the expressive ability of geometric features; (b) Intra-Graph Convolution: based on the sampling results, the local neighborhood graph is constructed and feature integration is performed to fuse cross-layer sampled multi-scale features, generating $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$.
@@
[…]
@@
Intra-Graph Convolution: The structure of the Intra-Graph Convolution is shown in Figure 4(b). It constructs a local neighborhood graph using the cross-layer sampled features $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$ output by the Cross-Layer Feature Point Sample and performs feature integration to generate $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$. It consists of two branches: the edge feature branch and the node feature branch, which are used to generate edge features and node features of the local neighborhood graph, respectively.
In the node feature branch, $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$ is first reconstructed into a 2D matrix $\mathcal{F}^{{}^{\prime}}_{g}\in\mathbb{R}^{N\times K\times C}$. After reconstruction, each row of $\mathcal{F}^{{}^{\prime}}_{g}$ corresponds to the neighborhood of a deep feature point, and the features within each row represent the shallow features in that neighborhood. Then, 2D convolution is used to extract features and generate node features $\mathcal{F}_{n}$. It should be noted that we use a 2D convolution with a kernel size of 1×31$\times$ 31 × 3 for feature extraction, which specifically aggregates features only within the same row. This design ensures that the neighborhoods of each deep feature point remain mutually independent, thereby avoiding mutual interference between node features across different neighborhoods.
In the edge feature branch, the reconstructed $\mathcal{F}^{{}^{\prime}}_{g}\in\mathbb{R}^{N\times K\times C}$ is matrix-multiplied with two learnable weight matrices $\mathcal{W}_{1}\in\mathbb{R}^{N\times C\times K}$ and $\mathcal{W}_{2}\in\mathbb{R}^{N\times K\times 1}$ to obtain edge point features $\mathcal{F}_{e}\in\mathbb{R}^{N\times K\times 1}$, which represent the correlation between the shallow feature points in each neighborhood and their corresponding deep feature points.
Finally, the node features and edge features are fused using element-wise multiplication and concatenated with $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$, and feature integration is performed through an MLP to generate ¯$\overline{\mathcal{F}}\in\mathbb{R}^{N\times C}$:
¯$\overline{\mathcal{F}}=MLP\left(Cat\left(Reshape\left(\mathcal{F}_{g}^{\prime}@\mathcal{W}_{1}@\mathcal{W}_{2}\right)\odot\mathcal{F}_{n},\ \mathcal{F}_{g}\right)\right)$ (9)
where $MLP(\cdot)$ denotes the Multi-Layer Perceptron; @@@ denotes matrix multiplication; $\mathcal{F}_{n}$ represents the node features; $Cat(\cdot)$ denotes the Concatenate operation; $Reshape(\cdot)$ denotes the Reshape operation.
@@ 3.5 Loss Function
Given the feature point pairs $[p_{i}^{src},f_{i}^{src}]\in[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $[p_{j}^{dst},f_{j}^{dst}]\in[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$, classify the paired samples into positive samples $\mathcal{Q}_{+}$, neutral samples $\mathcal{Q}_{o}$, and negative samples $\mathcal{Q}_{-}$ based on the distance between $p_{i}^{src}$ and $p_{j}^{dst}$, , following the rules of (10) to (12).
@@
Coarse Pairing Loss & Pairing Loss: We use the InfoNCE Loss [41] to supervise $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$ and $[\hat{\mathcal{P}},\hat{\mathcal{F}}]$, ensuring that the features of the source point cloud and the target point cloud are similar:
$\mathcal{L}_{c}=-\log\left(\frac{\sum_{\mathcal{Q}_{+}}\exp\left(\hat{f}_{i}^{src}\odot\hat{f}_{j}^{dst}/\alpha\right)}{\sum_{Q}\exp\left(\hat{f}_{i}^{src}\odot\hat{f}_{j}^{dst}/\alpha\right)}\right)\$ (13)
@@
$\mathcal{L}_{q}=\frac{1}{|\mathcal{Q}_{+}\cup\mathcal{Q}_{o}|}\sum_{\mathcal{Q}_{+}\cup\mathcal{Q}_{o}}\sigma^{p}_{i,j}\left\|\frac{q}{\|q\|}p_{i}^{src}\left(\frac{q}{\|q\|}\right)^{-1}+\tilde{t}-p_{j}^{dst}\right\|_{2}$ (16)
@@
Dynamic Segmentation Loss & Loop-Closure Detection Loss: We use cross-entropy to train the Dynamic Segmentation Head and the Loop-Closure Head. It is important to note that during the Loop-Closure training stage, we freeze the weights of the Encoder and Decoder components and only train the Loop-Closure Detection Head.
@@ 4.1 Implementation Details
To comprehensively evaluate the performance of ADA-DPM, we conducted experiments on the SemanticKITTI [42], KITTI-360 [43], and MulRan [44] datasets. Among them, sequences 00-05 of SemanticKITTI, sequences 00, 01, 03-06 of KITTI-360, as well as the Sejong sequence and DCC sequence of MulRan are used as the training sequences; sequences 05-10 of SemanticKITTI, sequences 07-10 and 18 of KITTI-360, along with sequences KAIST 03 (K03) and Riverside02 (R02) of MulRan are used as the test sequences. We adopted the standard Absolute Pose Error (APE) metric used in the KITTI-360 benchmark to quantitatively assess the global accuracy of the predicted trajectory.
@@
Training: We use the AdamW optimizer[45] with an initial learning rate of $lr=10^{-3}$, weight decay $wd=10^{-4}$, combined with a cosine learning rate scheduler[46]. The training is conducted on 3 RTX 3090 GPUs. For all tasks, the network is trained for 12 epochs, and the model from the final epoch is used for evaluation. Since only SemanticKITTI provides instance-level annotations for point clouds, we train the Dynamic Segmentation Head exclusively on SemanticKITTI. For training on the KITTI-360 and MulRan datasets, we utilize the pre-trained weights of the Dynamic Segmentation Head obtained from SemanticKITTI and freeze these weights (i.e., they do not participate in training).
Figure 6: Trajectory estimated result on different datasets.The blue line represents the predicted trajectory; the red line represents the ground-truth trajectory; the red pentagram represents the starting point; one grid represents 200m.
Table 1: Trajectories estimated results ($APE\downarrow$) on the test sequences. The bold indicates the best result, and the underline indicates the second-best result. GeoTransformer is a point cloud registration method, so for GeoTransformer, we perform end-to-end registration to complete the trajectory evaluation.

### v3

@@ 3.3 ADA-DPM for Point Cloud Registration
Importance Scoring Head: The Importance Scoring Head adopts an architecture of stacked MLP and KAN Linear Layer [24], as shown in Figure 3(b). The Importance Scoring Head treats the preliminary estimation of the rigid transformation between $\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$ as a regression task. By decoding $\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$, it predicts the rigid transformation between them and scores the matching situation as $\sigma_{p}\in\mathbb{R}^{N\times 1}$. Recently, Liu et al. [24] proposed a learnable KAN based on the Kolmogorov-Arnold theory. Unlike traditional approaches that rely on discontinuous activation functions, it learns the combination coefficients of continuous functions rather than the weights of node values, demonstrating strong performance in solving mathematical models. Therefore, during decoding, considering that the rigid transformation between the source and target point clouds satisfies strict mathematical constraints, we use the KAN Linear Layer as the decoder (its mathematical representation[wang2025s2kan] is shown Equations (4) and (5)).
@@
where $\Phi(\cdot)$ is the activation function matrix of the KAN layer; $\phi_{(i,j,k)}(\cdot)$ represents the learnable continuous activation function connecting the $k$-th neuron in the $i$-th layer and the $j$-th neuron in the $(i+1)$-th layer; $\mathcal{KAN}(\cdot)$ denotes a KAN Linear Layer.
It should be noted that directly predicting the rigid transformation $\mathcal{T}\in SE(3)$ between $\tilde{\mathcal{P}}_{src}$ and $\tilde{\mathcal{P}}_{dst}$ requires 12 variables to be predicted. To reduce the optimization difficulty, we adopt a simplification strategy: only predict the rotation transformation, which is represented using the quaternion $q=[s,(x,y,z)]$ (where $s$ denotes the real part of $q$, and $x,y,z$ are the imaginary parts of $q$); the translation transformation can be calculated using Equation (6). Specifically once $q$ is successfully predicted, we use $q$ to align the $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$. At this point, the difference in the centroid coordinates of the two aligned sets corresponds to the translation vector. In addition, by supervising $q$, implicit supervision can be imposed on the feature extraction network, enabling it to learn to extract feature points that contribute more to the point cloud registration task.
@@
To ensure that the Importance Scoring Head can properly score the matching of feature points, during the training phase, we specifically design the importance scoring loss $\mathcal{L}_{s}$ to supervise the prediction of $q$ and $\sigma_{p}$. $\mathcal{L}_{s}$ consists of the global reconstruction loss weighted by importance score $\mathcal{L}_{q}$ and the scoring loss $\mathcal{L}_{\sigma}$, where $\mathcal{L}_{q}$ is used to supervise the prediction of $q$, and $\mathcal{L}_{\sigma}$ is used to supervise the prediction of $\sigma_{p}$. Specifically, $\mathcal{L}_{q}$ is represented by the global reconstruction loss weighted by $\sigma_{p}$; $\mathcal{L}_{\sigma}$ is represented by cross-entropy with labels all being 1. Since the network weights are optimized in the direction that reduces $\mathcal{L}_{s}$ during training, for correctly paired point pairs, $\mathcal{L}_{q}$ can be reduced by predicting the correct $q$; for incorrectly paired point pairs, $\mathcal{L}_{q}$ needs to be reduced by predicting a smaller $\sigma_{p}$. Meanwhile, to prevent the model from reducing $\mathcal{L}_{q}$ by predicting small (or even all-zero) $\sigma_{p}$ for all point pairs, $\mathcal{L}_{\sigma}$ calculates the loss with a vector of all ones. This ensures that for correctly paired point pairs, the model predicts $\sigma_{p}$ as close to 1 as possible, and $\mathcal{L}_{q}$ can only be reduced by predicting the correct global rigid transformation, thus avoiding the trivial solution where $\sigma_{p}$ is extremely small or all zero in this case. At this point, $\sigma_{p}$ can be used to represent the contribution of point pairs to registration.
@@ 3.4 ADA-DPM for Loop-Closure Detection
For the Encoder part, we adopt the structure of PointNeXt [45] combined with CLI-GCN. Existing point cloud feature extraction methods (such as PointNet++ [34], VoxelNet [47], and PointNeXt [45] used in this paper) aggregate local features through downsampling. As a result, although the finally extracted compressed feature descriptors have rich semantic features, they suffer from issues such as a single receptive field and low feature resolution. For points in overlapping regions, changes in the size of their neighborhood structures do not affect their similarity; however, for points at the edges of overlapping regions, an increase in the receptive field will cause a sharp decrease in the similarity of their neighborhood structures, which may lead to incorrect loop-closure predictions and introduce errors into robot localization and map construction. To make up for this defect of PointNeXt, we construct the CLI-GCN module, which constructs neighborhood graph structures across feature layers to capture geometric features at different scales and emphasizes key geometric features, thereby enhancing the representation capability of the geometric features.
The CLI-GCN structure is shown in Figure 4, mainly consisting of two parts: (a) Cross-Layer Feature Point Sample: this part searches for neighborhood points for deep feature points within the shallow feature point set (performing neighborhood sampling based on distance), constructing multi-scale neighborhood correlations between deep feature points and shallow associated feature points to enhance the expressive ability of geometric features; (b) Intra-Graph Convolution: based on the sampling results, the local neighborhood graph is constructed and feature integration is performed to fuse cross-layer sampled multi-scale features, generating $\overline{\mathcal{P}},\overline{\mathcal{F}}]$.
@@
In addition, to avoid selecting feature points that overlap with deep feature points in 3D space or are extremely close to them, we set upper and lower distance thresholds ($0.5m$ and $1.0m$ in this paper), prioritizing the selection of shallow feature points within the distance threshold range. When the number of shallow feature points within the threshold range is less than $K$, supplementation is performed from feature points with distances greater than the upper threshold. Figure 5 shows the visualization results of deep feature points and feature points of the sampled local neighborhood graph. In Figure 5, blue points represent deep feature points, and orange points represent sampled shallow feature points. It can be seen that the blue feature points express more global structures, while the feature points of the local neighborhood graph obtained after sampling retain more local details, containing multi-scale geometric features of the point cloud.
@@
Intra-Graph Convolution: The structure of the Intra-Graph Convolution is shown in Figure 4(b). It constructs a local neighborhood graph using the cross-layer sampled features $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$ output by the Cross-Layer Feature Point Sample and performs feature integration to generate $\overline{\mathcal{P}},\overline{\mathcal{F}}]$. It consists of two branches: the edge feature branch and the node feature branch, which are used to generate edge features and node features of the local neighborhood graph, respectively.
In the node feature branch, $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$ is first reconstructed into a 2D matrix $\mathcal{F}^{{}^{\prime}}_{g}\in\mathbb{R}^{N\times K\times C}$. After reconstruction, each row of $\mathcal{F}^{{}^{\prime}}_{g}$ corresponds to the neighborhood of a deep feature point, and the features within each row represent the shallow features in that neighborhood. Then, 2D convolution is used to extract features and generate node features $\mathcal{F}_{n}$. It should be noted that we use a 2D convolution with a kernel size of $1\times 3$ for feature extraction, which specifically aggregates features only within the same row. This design ensures that the neighborhoods of each deep feature point remain mutually independent, thereby avoiding mutual interference between node features across different neighborhoods.
In the edge feature branch, the reconstructed $\mathcal{F}^{{}^{\prime}}_{g}\in\mathbb{R}^{N\times K\times C}$ is matrix-multiplied with two learnable weight matrices $\mathcal{W}_{1}\in\mathbb{R}^{N\times C\times K}$ and $\mathcal{W}_{2}\in\mathbb{R}^{N\times K\times 1}$ to obtain edge features $\mathcal{F}_{e}\in\mathbb{R}^{N\times K\times 1}$, which represent the correlation between the shallow feature points in each neighborhood and their corresponding deep feature points.
Finally, the node features and edge features are fused using element-wise multiplication (highlighting key node features with strong correlations) and concatenated with $\mathcal{F}_{g}\in\mathbb{R}^{(N\times K)\times C}$, and feature integration is performed through an MLP to generate $\overline{\mathcal{F}}\in\mathbb{R}^{N\times C}$:
$\overline{\mathcal{F}}=MLP\left(Cat\left(Reshape\left(\mathcal{F}_{g}^{\prime}@\mathcal{W}_{1}@\mathcal{W}_{2}\right)\odot\mathcal{F}_{n},\ \mathcal{F}_{g}\right)\right)$ (7)
where $MLP(\cdot)$ denotes the Multi-Layer Perceptron; @@ denotes matrix multiplication; $\mathcal{F}_{n}$ represents the node features; $Cat(\cdot)$ denotes the Concatenate operation; $Reshape(\cdot)$ denotes the Reshape operation.
@@ 3.5 Loss Function
Given the feature point pairs $p_{i}^{src},f_{i}^{src}]\in[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $p_{j}^{dst},f_{j}^{dst}]\in[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$, classify the paired samples into positive samples $\mathcal{Q}_{+}$, neutral samples $\mathcal{Q}_{o}$, and negative samples $\mathcal{Q}_{-}$ based on the distance between $p_{i}^{src}$ and $p_{j}^{dst}$, , following the rules of (8) to (10).
@@
Coarse Pairing Loss & Pairing Loss: We use the InfoNCE Loss [48] to supervise $\overline{\mathcal{P}},\overline{\mathcal{F}}]$ and $\hat{\mathcal{P}}$$\hat{\mathcal{F}}$], ensuring that the features of the source point cloud and the target point cloud are similar:
$\mathcal{L}_{c}=-\log\left(\frac{\sum_{\mathcal{Q}_{+}}\exp\left(\hat{f}_{i}^{src}\odot\hat{f}_{j}^{dst}/\alpha\right)}{\sum_{Q}\exp\left(\hat{f}_{i}^{src}\odot\hat{f}_{j}^{dst}/\alpha\right)}\right)\\$ (11)
@@
$\begin{split}\mathcal{L}_{q}&=\frac{1}{|\mathcal{Q}_{+}\cup\mathcal{Q}_{o}|}\\ &\sum_{\mathcal{Q}_{+}\cup\mathcal{Q}_{o}}\sigma^{p}_{i,j}\Big\|\frac{q}{\|q\|}p_{i}^{src}\left(\frac{q}{\|q\|}\right)^{-1}\quad+\tilde{t}-p_{j}^{dst}\Big\|_{2}\end{split}$ (14)
@@
Dynamic Segmentation Loss & Loop-Closure Detection Loss: We use cross-entropy to train the Dynamic Segmentation Head and the Loop-Closure Head. During the training of the Dynamic Segmentation Head, feature points corresponding to moving objects such as cars, pedestrians, and bicycles (including potentially movable objects like parked cars at road signs) are treated as dynamic points, while the remaining feature points are treated as static points. It is important to note that during the Loop-Closure training stage, we freeze the weights of the Encoder and Decoder components and only train the Loop-Closure Detection Head.
@@ 4.1 Implementation Details
Data Augmentation: (1) Random Sampling:For each input frame of point cloud data, randomly sample 16,384 points. If the number of input points is less than 16,384, pad with zeros to reach this count. (2) Random Occlusion [19]: Randomly mask 131–3 regions (removing all points within these regions) within an angular range of $10^{\circ}–150^{\circ}$ and a distance range of 2m10m2m–10m. (3) Random Rigid Transformation: Apply random rigid transformations to the input point cloud. Specifically, the rotation transformation is sampled from a uniform distribution with the mean of 0 and the standard deviation of $\pi$; the translation transformation is sampled from a uniform distribution with the mean of 0 and the standard deviation of 3. (4) Random Local Map Registration: Randomly sample 0n0–n frames of point clouds based on their distance from the current frame to construct a local map. Considering that in practical SLAM tasks, registration is usually made between the point cloud of the current frame and the local map built from historical frames, which expands over time. During training, we set $n$ to increase gradually as the number of training epochs increases. $n$ starts at 2, doubles every 3 epochs, and has a maximum value of 16.
Training: We use the AdamW optimizer[49] with an initial learning rate of $lr=10^{-3}$, weight decay $wd=10^{-4}$, combined with a cosine learning rate scheduler[50]. The training is conducted on 3 RTX 3090 GPUs. For all tasks, the network is trained for 12 epochs, and the model from the final epoch is used for evaluation.
@@
Modality Method 06 07 08 09 10
Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot. Trans. Rot.
LiDAR LOAM[15] 0.65 - 0.63 - 1.12 - 0.77 - 0.79 -
LO-Net[51] - - 0.56 0.45 1.08 0.43 0.77 0.38 0.92 0.41
ISC-LOAM[16] 0.76 0.41 0.56 0.43 1.20 0.50 1.40 0.59 1.87 0.62
SC-LeGO-LOAM[17] 2.54 1.15 2.48 1.78 2.30 1.24 5.37 2.78 10.50 3.79
F-LOAM[18] 0.84 0.33 0.88 0.62 0.87 0.33 1.03 0.32 1.20 0.29
LiODOM[52] 0.83 0.29 0.88 0.61 0.86 0.33 1.03 0.32 1.20 0.29
LiLO[53] 0.54 0.32 0.60 0.61 1.07 0.41 0.63 0.32 0.99 0.33
DeepPointMap[19] 0.77 - 0.44 - 1.09 - 0.95 - 0.69 -
Camera VISO2[54] 0.79 0.51 1.46 1.13 1.62 0.66 0.84 0.64 1.29 0.64
ORB-SLAM2[10] 0.89 0.27 0.89 0.50 1.03 0.31 0.86 0.25 0.62 0.29
Vins-Fusion[11] 1.35 0.71 1.21 0.90 1.83 0.72 1.82 0.53 2.64 1.01
OV2-SLAM[12] 1.13 0.28 1.03 0.57 1.11 0.31 0.96 0.20 0.52 0.18
SOFT2[13] 0.60 0.23 0.45 0.29 0.91 0.26 0.75 0.22 0.74 0.24
HVL-SLAM[4] 0.54 0.22 0.53 0.29 1.02 0.29 0.64 0.27 0.70 0.28
LiDAR&Camera DEMO[55] 0.96 - 1.16 - 1.24 - 1.17 - 1.14 -
DVL-SLAM[56] 0.92 - 1.26 - 1.32 - 0.66 - 0.70 -
CR-LDSO[57] 0.51 - 1.01 - 1.14 - 0.56 - 0.60 -
LiDAR ours 0.34 0.29 0.35 0.28 0.86 0.34 0.79 0.31 0.60 0.28
Figure 6: Trajectory estimation result on different datasets.The blue line represents the predicted trajectory; the red line represents the ground-truth trajectory; the red pentagram represents the starting point; one grid represents 200m.
4.2 Compared with Others
[…]
Trajectory Estimation: As shown in Figure 6, ADA-DPM achieves high-precision and globally consistent map reconstruction on all selected datasets. In the figure, the red line represents the ground-truth trajectory, the blue line represents the trajectory predicted by ADA-DPM, and the red pentagram marks the starting point of the trajectory. In particular, for long-range localization and mapping tasks like KITTI-360 (where KITTI-360: 09 contains point clouds over 10,000 meters), ADA-DPM still demonstrates favorable performance.
Table 2: Localization accuracy ($APE\downarrow$) on the test sequences. The bold indicates the best result, and the underline indicates the second-best result. GeoTransformer is a point cloud registration method, so for GeoTransformer, we perform end-to-end registration to complete the trajectory evaluation.